# BANG: Billion-Scale Approximate Nearest Neighbour Search using a Single GPU

### v2

@@ VII Results
In this section, we comprehensively compare the performance and solution quality of BANG with various state-of-the-art methods across different datasets detailed in Section VI. We present the results on 1B datasets followed by 100M datasets. Further, we also provide the cost analysis for BANG as supplementary material [2].
VII-A Performance on Billion-scale Datasets
Figure  5 compares BANG against GGNN and FAISS on billion-size datasets. BANG outperforms GGNN and FAISS in throughput for compatible recalls at 10-recall@10 across all three datasets. Overall, BANG achieves outstanding performance with 40$\times$-200$\times$ throughput for a high recall of 0.9.
@@ VII-A Performance on Billion-scale Datasets
For optimal performance in FAISS, we configure the index values to be large enough to utilise the available GPU memory to the maximum extent possible. However, FAISS reports relatively lower recall due to compression losses. The number of distance calculations exponentially increases with recall, as observed in [33], which negatively impacts its throughput. Further, in general, the distance calculations in FAISS are higher in order of magnitude than in proximity graph methods.
@@
For GGNN on a single GPU, data transfer overhead from CPU to GPU inlined with the search impeded the throughput. The memory transfer overhead was addressed by the original GGNN implementation by utilising eight GPUs to process the shards in parallel. For instance, after utilizing the 80 GB of GPU global memory, SIFT1B requires an additional 143 GB of data to be transferred from the CPU to the GPU during the search since the dataset and the graph are each 128 GB and 95 GB (total 223 GB), respectively. Transferring this additional data over the PCIe bus @ 32GB/s will take close to 5 seconds.
@@
BANG search runs on GPU but does not require the graph to reside on the GPU. It further eliminates the need to transfer the base dataset to GPU by employing compressed vectors (instead of base vectors) present on the GPU. Furthermore, the reason for the high throughput of BANG is attributed to the massive parallelism achieved on the GPU (Section IV). The re-ranking step, coupled with the integration of Bloom filters to filter processed nodes (prevents duplicate entries in the worklist and allows more deserving nodes to be included in the worklist), significantly contributes to achieving a high recall for BANG.
@@ VII-B Performance on 100M Datasets
Figure 6 shows the comparison of throughput and recall on 100M datasets. As the data structures (i.e., graph and data points) for 100M datasets fit within GPU memory, we report the performance of the two variants of BANG (Section V). The BANG In-memory and BANG Exact-distance variants achieve 2$\times$ and 3$\times$ higher throughput compared to the base variant of the BANG since there is no CPU-GPU data transfer overhead in the BANG variants. Interestingly, BANG Exact-distance variant outperforms GGNN. We might expect the same performance from the BANG Exact-distance variant as GGNN since it uses only GPU-resident data structures as done in GGNN and does not use compressed vectors for distance computations to query points. However, the main difference between the two is the type of underlying proximity graph. The k-NN graph structure of GGNN requires more hops or iterations for the search to terminate. In contrast, in BANG Exact-distance, the Vamana graph contains long-range edges, reducing the hops, enabling faster convergence with fewer distance computations. Note that, compared to BANG and FAISS, for GGNN, the performance improvement on 100M datasets over 1B datasets is significant since the CPU to GPU data transfer overhead, which is present in the 1B datasets, is not present here.
Figure 7: Variation in throughput and recall with respect to $m$ on SIFT1B.
VII-C Effect of Varying Compression Ratio
@@
PQ compression techniques invariably introduce a loss in the recall. To empirically determine the impact of the compression operation on BANG’s recall, we ran BANG search on the SIFT1B dataset and varied $m$ (i.e. the number of subspaces) during each run while keeping the other configurations the same as in Section VI. The results are shown in Figure 7.
@@
While we expect the recall to decrease with high compression (i.e. a lower value of $m$), the result shows no noticeable change in throughput or recall values even when $m$ is decreased from 74 to 32 (representing compression ratios of 0.57 and 0.25, respectively). We observed that recall does not drop until a compression ratio of 0.25; below this value, recall begins to decrease. Thus, this behaviour allows BANG to operate efficiently on GPUs with limited device memory without compromising recall or throughput. For example, we expect BANG search to deliver comparable recall and throughput on an A100 GPU configured with either 40 GB or 80 GB of device memory. This translates to significant energy and cost savings.
@@
One would expect the throughput to increase with the reduction in compression ratio because it reduces the number of distance computations (requiring fewer PQ distance lookups as the number of subspaces reduces) in the distance calculation kernel. Interestingly, we do not notice this pattern in the results. The inaccuracies in the distances calculated using compressed vectors increase as $m$ decreases. Hence, BANG search path needs more hops and detours to converge. The additional time spent in performing more search iterations outweighs the gains from the computations saved on distance calculations, resulting in no significant throughput improvement.
@@
TABLE III: Variation of $\lambda$ with respect to search parameter $t$, i.e. the size of the worklist. (Smaller $\lambda$ values are better.)
@@
VII-D Ablation Study
We varied $t$ from 20 through 180 and measured $\lambda$ and 10-recall@10 for the SIFT1B dataset with $m=64$ chunks. Table III presents the results. Note that recall increases with $t$. So, the purpose of considering larger values of $t$ is to increase the recall (as it enables exploring more search paths). As we can observe from Table III, $\lambda$ decreases with an increase in $t$. A lower value of $\lambda$ indicates faster convergence with fewer iterations, confirming the efficacy of the recall improvement strategies for the graph index used.
Furthermore, in each iteration, multiple queries are processed in parallel, each by a separate thread block. The number of queries processed in the first iteration equals the batch size. In our phased-execution strategy, each query takes one step (a phase) at a time, along with all the other queries in the batch. The queries that are converged to termination are excluded from subsequent iterations. The total number of iterations ($I$) for the greedy search to complete (i.e., for all queries in a batch to converge) is crucial for throughput; fewer iterations result in higher throughput. For a particular $t$ value, we observed that on average, 95% of queries in a batch were completed in $1.1\times t$ iterations. This demonstrates that most queries take close to the optimal ($t$) number of iterations to complete. Hence, the parallelization of BANG does not require sophisticated and heavy-weight thread scheduling mechanisms such as work-stealing wherein a thread-block which has finished processing its assigned query $q_{i}$ can partake in processing unprocessed candidate nodes in the worklist $\mathcal{L}_{j}$ of another query $q_{j}$.
@@ VIII Related Work
ANNS on CPU. Traditionally, ANNS has utilized Tree-based methods (like KD-trees [8, 14] and cover trees) and followed a branch-and-bound approach for search navigation. Hashing-based methods are also explored for ANNS. Locality-Sensitive Hash (LSH) [3, 4] depends on developing hash functions that can generate higher collision probabilities for nearby points than those far apart. Both these approaches face scalability challenges due to the growing dimensions and size.
@@
NSW [32] is an early graph-based ANNS method that employs the Delaunay Graph (DG) [17] for identifying node neighbours, ensuring near-full connectivity but introducing a high-degree search space with increasing dataset dimensionality. The graph includes short-range edges for higher accuracy and long-range edges for improved search efficiency, but this results in an imbalance in the graph’s degree due to the creation of “traffic hubs”. HNSW [31] addresses this by introducing a hierarchical structure to spread neighbours across levels and imposing an upper bound on node degree. However, scaling HNSW to high dimensions and large datasets continues to be challenging. NSG [18] improves graph-based methods’ efficiency and scalability, introducing the Monotonic Relative Neighborhood Graph (MRNG) for reducing index size and search path length while also scaling to billion-scale datasets. DiskANN [25] indexes billion-point datasets in hundreds of dimensions through the Vamana graph on commodity hardware (just 64GB RAM and an inexpensive solid-state drive), leveraging NSG and NSW strengths. It introduces a tunable parameter $\alpha$ for a graph degree-diameter trade-off during construction. It proposes compression schemes to reduce memory consumption and computation costs during the search (see Section II-B). SPANN [9] utilizes the inverted index methodology. Unlike other inverted index techniques that use compression techniques (that have some invariable losses) to reduce memory footprint, SPANN uses SSD to store a portion of the index and yet outperforms its predecessors by greatly reducing the number of disk accesses. HM-ANN [38] uses memory technologies like Optane PMM to expand the main memory size to fit the graph index and thus avoid compression. It extends popular graph-based techniques like HNSW and NSG to run on the heterogeneous memory architecture with billion-scale datasets. The survey by Manohar et al. [16] extensively studies and compares popular graph-based ANNS algorithms, addressing scalability challenges and proposing parallelization strategies on CPU.
@@
ANNS on GPU. Due to the complexity and throughput requirement of ANNS, offloading these calculations to massively parallel accelerators like GPU has gained considerable attention. FAISS [27] utilizes Quantization techniques for dimensionality reduction, employing an inverted index as the underlying data structure. FAISS achieves high throughputs at a billion scale, but with lower recall values. PQT [45] extends Product Quantization for better GPU performance but has drawbacks, such as longer encoding lengths than PQ-code and high encoding errors. V-PQT [10] addresses this with Vector and Product Quantization Tree, and Parallel-IVFADC [41] proposes a distributed memory version of FAISS for hybrid CPU-GPU systems scaling up to 256 nodes.
@@
SONG [49], a graph-based ANNS implementation on GPU, offers a notable speedup over state-of-the-art CPU counterparts and significant improvements over FAISS through efficient GPU-centric strategies for ANNS. GANNS [47], built on this foundation, explores enhanced parallelism and occupancy through GPU-friendly data structures and an NSW-based proximity graph construction scheme. In a related effort, TSDG [43] proposes a two-stage graph diversification approach for graph construction and GPU-friendly search procedures catering to various batch query scales. CAGRA [35], a recent graph-based ANNS implementation, leverages modern GPU capabilities (e.g., warp splitting) for substantial performance gains over existing GPU ANNS methods at a million scale. However, SONG, TSDG, GANNS and CAGRA face limitations in scaling to billion-scale datasets due to the necessity of the entire graph index to reside on the GPU.
@@
Graph-based GPU Nearest Neighbor (GGNN) [21] search is a recent implementation that outperforms SONG. It divides the large input dataset into smaller shards that can fit into GPU memory for performing ANNS on GPU, with the final results merged on the CPU. GGNN can run on a single GPU at a billion scale but with low throughput. To achieve very high throughputs ($>$ 100,000) at high recalls ($>$ 0.95), it utilizes eight GPUs for billion-scale ANNS.
@@
ANNS on other Accelerators. Researchers have explored custom hardware, such as FPGAs, for ANNS acceleration, focusing on quantization-based [1, 48] and graph-based [36] approaches. However, these methods suffer from frequent data movement between CPU memory and device memory, leading to high energy consumption and lower throughputs. Addressing these limitations, vStore [30] proposes an in-storage computing technique for graph-based ANNS within SSDs, avoiding data movement overhead and achieving low search latency with high accuracy. On the other hand, TPU-KNN [11] focuses on realizing ANNS on TPUs. It uses an accurate accelerator performance model considering both memory and instruction bottlenecks.

### v4

@@ VII Results
In this section, we comprehensively compare the performance and solution quality of BANG with various state-of-the-art methods across different datasets detailed in Section VI. We present the results on 1B datasets followed by 100M datasets.
VII-A Performance on 1B Datasets
Figure 5 compares BANG against GGNN and FAISS on billion-size datasets. BANG outperforms GGNN and FAISS in throughput for compatible recalls at 10-recall@10 across all four datasets. On the SIFT-1B and Deep-1B datasets, BANG achieves an average (geomean) throughput improvement of 50.5$\times$ over the competing methods at a high 10-recall@10 of 0.95. Specifically, for a 10-recall@10 of 0.9, BANG achieves speedups of 55$\times$, 50$\times$, and 400$\times$ on the SIFT-1B, Deep-1B, and MSSPACEV-1B datasets, respectively. On the more challenging Text-to-Image-1B dataset, BANG and other methods exhibit lower recall values. Here, too, BANG outperforms the closest baseline by 25$\times$ at a 10-recall@10 of 0.8.
@@ VII-A Performance on 1B Datasets
To achieve optimal performance in FAISS-PQ, we configure the index values to be large enough to utilise the available GPU memory to the maximum extent possible. However, FAISS-PQ fails to achieve higher recall values (compared to graph-based techniques) because of losses due to compression. The number of distance calculations exponentially increases with recall, as observed in [38], which negatively impacts its throughput. Furthermore, in general, the number of distance computations in FAISS is orders of magnitude higher than in proximity graph-based methods. FAISS-SQ achieves higher recall due to milder compression, as each dimension is encoded with a single byte. In contrast, FAISS-PQ encodes $M$ dimensions using one byte, resulting in stronger compression but lower recall. However, FAISS-SQ’s larger index size requires data swaps between the CPU and GPU, leading to lower throughput compared to FAISS-PQ.
@@
For GGNN on a single GPU with 1B dataset, data transfer overhead from CPU to GPU inlined with the search impedes the throughput. After utilizing the 80 GB of GPU global memory: the SIFT-1B dataset, for example, requires an additional 123 GB of data to be transferred from the CPU to the GPU during the search. Since the dataset and the graph are each 128 GB and 75 GB (total 203 GB), respectively. Transferring this additional 123 GB data over the PCIe bus @ 32 GB/s will take more than 3.8 seconds in practice. The original GGNN implementation addressed the memory transfer overhead by using eight GPUs to process the shards in parallel.
@@
BANG search runs on GPU but does not require the graph to reside on the GPU. It eliminates the need to transfer the base dataset to GPU by employing compressed vectors (instead of base vectors) present on the GPU. Furthermore, the reason for the high throughput of BANG is attributed to the massive parallelism achieved on the GPU (Section IV). The re-ranking step, coupled with the integration of Bloom filters to filter processed nodes (prevents duplicate entries in the worklist and allows more deserving nodes to be included in the worklist), significantly contributes to achieving a high recall for BANG.
Figure 6 presents the results of our experiments with 100-recall@100. While ANNS evaluations in the literature typically focus on 10-recall@10, real-world applications like recommender systems [58] often require fetching a larger number of nearest neighbours using ANNS, which are then filtered for final recommendations. To reflect this, we extend our evaluation to include 100-recall@100. The drop in QPS for BANG is due to the increased worklist length $t$ required to achieve the same recall as 10-recall@10, with the minimum value of $t$ being $k$, where $k$ is 100 for 100-recall@100. (The effect of varying worklist length on recall and throughput is detailed in Section VII-G). There is no significant throughput deterioration in FAISS, as the same number of clusters and entries within the posting lists are compared regardless of the recall parameter. GGNN, however, exhibits low 100-recall@100 values ($<0.1$), so we omit its plots in Figure 6.
TABLE III: Breakdown of the total ANNS execution time for 10-recall@10 = 0.9 and query batch size = 10,000.
Activity BANG GGNN FAISS-SQ
SIFT-1B Deep-1B SIFT-1B Deep-1B SIFT-1B Deep-1B
Time (ms) Time (ms) Time (ms) Time (ms) Time (ms) Time (ms)
GPU Processing 65.8 77.5 284.8 496.0 13.5 5.7
CPU Processing 15.9 20.6 0.1 0.1 14107.0 6963.9
Data Transfer (CPU $\leftrightarrow$ GPU) 15.4 16.4 5087.6 16303.9 9.6 9.9
Overall Processing 85.3 101.2 5372.5 16800 14130.1 6979.5
TABLE IV: Percentage breakdown of GPU time (from values in Table III), and memory usage of BANG kernels.
Kernel SIFT-1B Deep-1B
Time (%) Mem (MB) Time (%) Mem (MB)
Compute PQ Distance table 2 724 2 726
Calc neighbours’ PQ dist to query point 46 71300 49 71300
Sort neighbours and merge with Worklist 18 20 17 21
Filter out visited neighbours 20 3819 19 3819
Calc exact dist of all candidates (Re-rank) 4 158 4 448
Pre-fetch next candidate 10 10 9 18
@@ VII-B Performance on 100M Datasets
Figure 7 shows the comparison of throughput and recall on 100M datasets. As the data structures (i.e., graph and data points) for 100M datasets fit within GPU memory, we report the performance of the two variants of BANG (Section V). The BANG In-memory and BANG Exact-distance variants achieve 2$\times$ and 3$\times$ higher throughput compared to the base variant of BANG since there is no CPU-GPU data transfer overhead in the BANG variants. Interestingly, the BANG Exact-distance variant outperforms GGNN. We might expect the same performance from the BANG Exact-distance variant as GGNN, since it uses only GPU-resident data structures as in GGNN and does not use compressed vectors for distance computations to query points. However, the main difference between the two is the type of underlying proximity graph. The k-NN graph structure of GGNN requires more hops or iterations for the search to terminate than BANG Exact-distance. In contrast, in BANG Exact-distance, the Vamana graph contains long-range edges, reducing the hops, enabling faster convergence with fewer distance computations. Note that, compared to BANG and FAISS, for GGNN, the performance improvement on 100M datasets over 1B datasets is significant, since the CPU to GPU data transfer overhead, which is present in the 1B datasets, is not present here.
We compare with CAGRA on all 100M datasets except Text-to-Image, because the dataset (80 GB) and graph index (12.8 GB), which are required for its execution, exceed the 80 GB global memory capacity of the A100 GPU. We observe that CAGRA achieves nearly 1.5×1.5$\times$1.5 ×2×2$\times$2 × higher throughput than BANG, although it performs 13×13$\times$13 × more distance computations than BANG. Furthermore, using the NVIDIA Nsight-Compute profiler (NCU) on the Deep-100M dataset, we measure nearly 7 billion floating-point operations for BANG and nearly 12 billion for CAGRA. However, CAGRA achieves a higher memory throughput of 1270 GB/s compared to 577 GB/s for BANG. It also has a small memory access overhead due to uncoalesced accesses of 3% compared to 40% for BANG, as measured by NCU using the metric derived__memory_l2_theoretical_sectors_global_excessive [44]. The higher performance of CAGRA is because it uses the highly optimised RAFT [42] primitives for bitonic sort/merge and a space-efficient hash table that fits in shared memory, reducing the access latency. As we target billion-scale datasets, our frequently accessed bloom filter data structure is too large to fit in shared memory; this results in frequent global memory transactions, which impedes performance.
VII-C Search Latency and Memory Usage
To gain insight into the search latency, we study the time spent by different ANNS methods on (i) CPU execution, (ii) GPU execution, (iii) CPU-GPU data transfer, and (iv) total search time. Table III presents the results for BANG and baseline methods on two 1B datasets at a 10-recall@10 value of 0.9. For baselines, the overall processing time (the last row) is the sum of the first three components ((i)-(iii)). However, for BANG, due to its CPU-GPU load balancing with prefetching and asynchronous transfers, the total time is reduced by approx. $12\%$ on both datasets compared to its unoptimised implementation. Further, the table confirms that GGNN suffers from high transfer overheads due to frequent shard swaps. FAISS-SQ performs only coarse quantization on the GPU and actual search on the CPU, resulting in low GPU processing time. FAISS-PQ variant does not reach the target 10-recall@10 of 0.9 and is thus excluded from the table. For completeness, FAISS-PQ’s execution times for 10-recall@10 of 0.7 on SIFT-1B and DEEP-1B are {4154, 45.9, 0.06, 4200} ms and {4302, 24.9, 0.07, 4327} ms, respectively, for (i) CPU, (ii) GPU, (iii) transfer, and (iv) total time. As these values indicate, unlike FAISS-SQ, the entire search in FAISS-PQ is performed on the GPU.
Furthermore, we capture the peak GPU memory usage (in GB) of each method, out of the 80 GB available on the A100 GPU. On SIFT-1B, it is 76.0 (BANG), 78.8 (GGNN), 44.6 (FAISS-PQ), and 2.6 (FAISS-SQ); on Deep-1B, it is 76.3, 68.8, 45.9, and 2.5, respectively. The memory usage is dominated by compressed vectors for BANG, posting lists for GGNN, and graph index shards for FAISS-PQ. Thanks to PQ compression, BANG can run on lower-memory GPUs (e.g., 40GB A100 GPU) while maintaining comparable performance (Section VII-E). FAISS-SQ has the least GPU-memory usage, as only centroids reside on the GPU, with search executed on the CPU. For BANG, we further measure the distribution of GPU time and memory usage for each kernel. Table IV presents the results. The majority of the time and memory is utilised by the key kernels involved in distance calculation and neighbour filtering, which is consistent with the analyses in Sections IV-D and IV-C.
VII-D Evaluation of Optimisations in BANG
Key optimisations in BANG include multi-kernel approach, prefetching candidate nodes and asynchronous memory transfers between the host and GPU, discussed in IV, IV-E, and Sections IV-B, respectively. We perform experiments on two popular 1B datasets at a baseline 10-recall@10 value of 0.9 to demonstrate the efficacy of the three important optimisations. To measure their impact, we disabled each optimisation individually and recorded the throughput. The performance improvements for SIFT-1B are 10.1%, 8%, and 6.2%, and for Deep-1B are 17.1%, 16.5%, and 13.8%, respectively. These results demonstrate that the proposed optimisations are significant contributors to BANG’s superior performance.
VII-E Effect of Varying Compression Ratio
Figure 8: Throughput (log-scale) vs. 10-recall@10 for BANG on SIFT-1B with varying compressed sizes; $m$ = number of vector subspaces employed in compression (see Section IV).
@@ VII-E Effect of Varying Compression Ratio
PQ compression techniques invariably introduce a loss in the recall. To empirically determine the impact of the compression operation on BANG’s recall, we ran BANG search on the SIFT-1B dataset and varied $m$ (i.e. the number of subspaces) during each run while keeping the other configurations the same as in Section VI. The results are shown in Figure 8.
@@
While we expect the recall to decrease with high compression (i.e. a lower value of $m$), the result shows no noticeable change in throughput or recall values even when $m$ is decreased from 74 to 32 (representing compression ratios of 0.57 and 0.25, respectively). We observed that recall does not drop until a compression ratio of 0.25; below this value, recall begins to decrease. Thus, this behaviour allows BANG to operate efficiently on GPUs with limited device memory without compromising recall or throughput. For example, we expect BANG search to deliver comparable recall and throughput on an A100 GPU configured with either 40 GB or 80 GB of device memory. This translates to significant power and cost savings.
@@
One would expect the throughput to increase with the reduction in compression ratio because it reduces the number of distance computations (requiring fewer PQ distance lookups as the number of subspaces reduces) in the distance calculation kernel. Interestingly, we do not notice this pattern in the results. The inaccuracies in the distances calculated using compressed vectors increase as $m$ decreases. Hence, BANG’s search path needs more hops and detours to converge. The additional time spent in performing more search iterations outweighs the gains from the computations saved on distance calculations, resulting in no significant throughput improvement.
VII-F Effect of Varying Query Batch Size
We analyse the impact of input query batch size on throughput by varying it from 1 to 20,000 using the MSSPACEV-1B dataset, which contains more than 10,000 queries. To include all the baseline algorithms for comparison, we configure the parameters to achieve a 10-recall@10 of $\sim$0.8. As shown in Figure 9, since BANG and the baselines assign one query per thread block, increasing the batch size allows greater GPU concurrency, improving QPS. Throughput increases rapidly for small batch sizes due to more queries running in parallel on all the available GPU cores but stabilises beyond a saturation point when all GPU cores are fully utilised. The saturation points for BANG, FAISS-PQ, and FAISS-SQ occur at batch sizes of approximately 4000, 1000, and 400, respectively. BANG ’s phased execution strategy enables effective parallelism at higher batch sizes.
For GGNN, we observe a relatively higher increase in throughput in relation to the increase in batch size compared to other methods. As discussed in Section VII-C, the data transfer time highly dominates the query processing time. Since the data transfer time remains constant and independent of batch size, increasing the batch size leads to a proportionally greater increase in the numerator (number of queries) compared to the denominator (search time), resulting in a relatively higher improvement in throughput for GGNN.
Figure 9: Throughput (log-scale) vs. query batch size on MSSPACEV-1B dataset. Batch size denotes the number of queries processed concurrently.
VII-G Effect of Varying Worklist Size ($t$)
@@
TABLE V: Variation of $\lambda$ with respect to search parameter $t$, i.e. the size of the worklist, on SIFT-1B dataset. (smaller $\lambda$ values lead to higher throughputs)
@@
We vary $t$ from 20 through 180 and measure $\lambda$ and 10-recall@10 for the SIFT-1B dataset with $m=64$ chunks. Table V presents the results. Note that recall increases with $t$. So, higher values of $t$ increase the recall by enabling the exploration of more search paths. As we can observe from Table V, $\lambda$ decreases with an increase in $t$. A lower value of $\lambda$ indicates faster convergence with fewer iterations, confirming the efficacy of the recall improvement strategies for the graph index used.
Furthermore, in each iteration, multiple queries are processed in parallel, each by a separate thread block. The number of queries processed in the first iteration equals the batch size. In our phased-execution strategy, each query takes one step (a phase) at a time, along with all the other queries in the batch. The queries that are converged to termination are excluded from subsequent iterations. The total number of iterations ($I$) for the greedy search to complete (i.e., for all queries in a batch to converge) is crucial for throughput; fewer iterations result in higher throughput. For a particular $t$ value, we observed that on average, 95% of queries in a batch are completed in $1.1\times t$ iterations. This demonstrates that most queries take close to the optimal ($t$) number of iterations to complete. Hence, the parallelisation of BANG does not require sophisticated and heavy-weight thread scheduling mechanisms.
@@ VIII Related Work
ANNS on CPU. Traditionally, ANNS has utilised Tree-based methods (like KD-trees [9, 15] and cover trees) and followed a branch-and-bound approach for search navigation. Hashing-based methods are also explored for ANNS. Locality-Sensitive Hash (LSH) [2, 3] depends on developing hash functions that can generate higher collision probabilities for nearby points than those far apart. Both these approaches face scalability challenges due to the growing dimensions and size.
@@
NSW [37] is an early graph-based ANNS method that employs the Delaunay Graph (DG) [20] for identifying node neighbours, ensuring near-full connectivity but introducing a high-degree search space with increasing dataset dimensionality. The graph includes short-range edges for higher accuracy and long-range edges for improved search efficiency, but this results in an imbalance in the graph’s degree due to the creation of traffic hubs. HNSW [36] addresses this by introducing a hierarchical structure to spread neighbours across levels and imposing an upper bound on node degree. However, scaling HNSW to high dimensions and large datasets continues to be challenging. NSG [21] improves graph-based methods’ efficiency and scalability, introducing the Monotonic Relative Neighborhood Graph (MRNG) for reducing index size and search path length while also scaling to billion-scale datasets. DiskANN [29] indexes billion-point datasets in hundreds of dimensions through the Vamana graph on commodity hardware (having a 64GB RAM and an inexpensive solid-state drive), leveraging the properties of NSG and NSW. It introduces a tunable parameter $\alpha$ for a graph degree-diameter trade-off during construction. It proposes compression schemes to reduce memory consumption and computation costs during the search (see Section II-B). SPANN [11] utilises the inverted index methodology. Unlike other inverted index techniques that use compression techniques (that have some invariable losses) to reduce memory footprint, SPANN uses SSD to store a portion of the index and yet outperforms its predecessors by greatly reducing the number of disk accesses. HM-ANN [51] uses memory technologies like Optane PMM to expand the main memory size to fit the graph index and thus avoid compression. It extends popular graph-based techniques like HNSW and NSG to run on the heterogeneous memory architecture with billion-scale datasets. The survey by Manohar et al. [18] extensively studies and compares popular graph-based ANNS algorithms, addressing scalability challenges and proposing parallelization strategies on CPU.
@@
ANNS on GPU. Due to the complexity and throughput requirement of ANNS, offloading these calculations to massively parallel accelerators like GPU has gained considerable attention. FAISS [31] utilises quantization techniques for dimensionality reduction. The underlying search structure of FAISS is the inverted-index (IVF) [40], where points are typically assigned to buckets using either single-level clustering or two-level clustering. FAISS achieves high throughputs on the billion scale but with low recall values. PQT [61] extends Product Quantization for better GPU performance but has drawbacks, such as longer encoding lengths than PQ-code and high encoding errors. V-PQT [12] addresses this with Vector and Product Quantization Tree, and Parallel-IVFADC [54] proposes a distributed memory version of FAISS for hybrid CPU-GPU systems scaling up to 256 nodes.
@@
SONG [67], a graph-based ANNS implementation on GPU, offers a notable speedup over state-of-the-art CPU counterparts and significant improvement over FAISS through efficient GPU-centric strategies for ANNS. GANNS [63], built on this foundation, explores enhanced parallelism and occupancy through GPU-friendly data structures and an NSW-based proximity graph construction scheme. In a related effort, TSDG [56] proposes a two-stage graph diversification approach for graph construction and GPU-friendly search procedures catering to various batch query scales. CAGRA [46], a recent graph-based ANNS implementation, leverages modern GPU capabilities (e.g., warp splitting) for substantial performance gains over existing GPU ANNS methods at a million scale. However, SONG, TSDG, GANNS and CAGRA face limitations in scaling to billion-scale datasets due to the necessity of the entire graph index to reside on the GPU.
@@
Graph-based GPU Nearest Neighbor (GGNN) [24] search is a recent implementation that outperforms SONG. It proposes a hierarchical $k$NN graph construction algorithm. It divides the large input dataset into smaller shards that can fit into GPU memory for performing ANNS on GPU, with the final results merged on the CPU. GGNN can run on a single GPU at a billion scale but with low throughput. To achieve very high throughputs ($>100,000$) at high recalls ($>0.95$), it utilises eight GPUs for billion-scale ANNS.
@@
ANNS on other Accelerators. Researchers have explored custom hardware, such as FPGAs, for ANNS acceleration, focusing on quantization-based [1, 64] and graph-based [48] approaches. However, these methods suffer from frequent data movement between CPU memory and device memory, leading to high energy consumption and lower throughputs. To address these limitations, vStore [35] proposes an in-storage computing technique for graph-based ANNS within SSDs, avoiding data movement overhead and achieving low search latency with high accuracy. On the other hand, TPU-KNN [13] focuses on realizing ANNS on TPUs. It uses an accurate accelerator performance model considering both memory and instruction bottlenecks.